OpenStudy (anonymous):

Solve the following system of equations. 2x + y = 3 x = 2y - 1

hartnn (hartnn):

plug in x=2y-1 in the first equation, 2x +y=3 you get 2 (2y-1)+y=3 can you solve this equation which is only in 'y' and find y ?

OpenStudy (anonymous):

y+5/6

hartnn (hartnn):

how ? 4y-2+y=3 5y=3+2 5y=5 y=1 got this ?

OpenStudy (anonymous):

i didnt distribute first i did something different ; how would i do this one Solve the following system of equations. 3x + 2y - 5 = 0 x = y + 10

hartnn (hartnn):

put x=y+10 in 3x+2y-5=0 so, 3(y+10) +y -5=0 distribute first.

OpenStudy (anonymous):

y=8.75

hartnn (hartnn):

3y+30+y-5=0 4y+25=0 y = -25/4 = - 6.25 how u got 8.75 ?

OpenStudy (anonymous):

i add 30+5 instead of subtracting

hartnn (hartnn):

k, now clear ?

OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3

hartnn (hartnn):

new question , new post plese ?